A man is now five times as old as his daughter and five years ago the product of their ages was 105. Find their present ages.
Answer by ikleyn(52871)

.
M = 5*D               (1)
(M-5)*(D-5) = 105     (2)


Substitute (1) into (2):

(5D-5)*(D-5) = 105

(D-1)*(D-5)  =  21

D^2 -D - 5D + 5 = 21

D^2 - 6D - 16 = 0

(D+2)*(D-8) = 0.


The positive root D= 8 is the only meaningful solution.


Answer.  Daughter is 8 years old.  Father is 40 years old.

Solved.
----------------
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.